# Robust Market Interventions

### v1

@@
Abstract.
A large differentiated oligopoly yields inefficient market equilibria. An authority with imprecise information about the primitives of the market aims to design tax/subsidy interventions that increase efficiency robustly—i.e., with high probability. We identify a condition on demand that guarantees the existence of such interventions, and we show how to construct them using noisy estimates of demand complementarities and substitutabilities across products. The analysis works by deriving a novel description of the incidence of market interventions in terms of spectral statistics of a Slutsky matrix. Our notion of recoverable structure ensures that parts of the spectrum that are useful for the design of interventions are statistically recoverable from noisy demand estimates.
Galeotti is at the London Business School, Golub is at Northwestern University, Goyal is at University of Cambridge and NYUAD, Talamàs is at IESE Business School, and Tamuz is at Caltech. We thank Yann Calvó-López and Yu-Chi Hsieh for excellent research assistance. This work was supported by the ERC grant 724356 (Galeotti), the Pershing Square Fund for Research on the Foundations of Human Behavior (Golub), the National Science Foundation (SES-1658940, SES-1629446, Golub & DMS-1944153, Tamuz), the Sloan Foundation (Tamuz) and the US-Israel Binational Science Foundation (2018397, Tamuz). We thank Bruno Pelligrino for valuable exchanges and kindly sharing data with us, and Volker Nocke for a helpful discussion on a related project at the 2023 AEA meetings. We thank Pol Antràs, Modibo Camara, Yeon–Koo Che, Kevin Chen, Jeff Ely, Matthew Gentzkow, Tim Gowers, Matthew O. Jackson, Jose Luis Moraga, Stephen Morris, Martin Pietz, Ran Spiegler, Xavier Vives, and Kuang Xu for helpful comments.
@@ 1. Introduction
The growing importance of firms with market power has led to a renewed interest in the study of imperfectly competitive markets (see, e.g., De Loecker, Eeckhout, and Unger (2020); Azar and Vives (2021); Pellegrino (2021); Ederer and Pellegrino (2021)). Markets with oligopolistic structures typically yield inefficient equilibrium outcomes, which can be improved through interventions. Intervention on one firm, however, creates spillovers on other firms due to their interrelated demands. An authority often has limited information on the structure of market demands. This leads us to develop a theory to determine when and how an authority with imprecise demand information can design interventions that increase welfare. Our approach combines classical welfare pass-through theory with statistical results on recoverable latent structures of large matrices to identify conditions on demand structure that ensure the robust achievement of welfare improvements even when many aspects of the demand system cannot be accurately estimated.
We study a multi-market oligopoly model—with general complementarities and substitutabilities across products—in which firms simultaneously set prices. The authority may be a marketplace operator (such as Alibaba or Amazon) or a government agency that regulates the market. This authority has access to a noisy signal about the market demand parameters—for instance, estimates produced from data on consumer behavior. We focus on intervention rules that prescribe a firm-specific tax/subsidy intervention for each possible signal realization. Our interest lies in the conditions under which intervention rules can increase total surplus robustly—with high probability, conditional on any true market state.
@@
What makes the problem challenging is that, in markets with numerous and changing goods, such as those hosted on large online platforms, the demand system is very high-dimensional, and realistic signals will leave substantial uncertainty about many aspects of the market’s structure; see Section 7.1 for a detailed discussion. Our main result is that if demand satisfies a property that we call “recoverable structure,” then there are feasible intervention rules that robustly increase the total surplus. Moreover, within a natural class of interventions—those which do not reduce consumer surplus [^1] This constraint is natural for a platform that could lose customers to other marketplaces or a government that can lose political support.—our feasible interventions achieve the largest gain in surplus that is possible for a given level of subsidy expenditure. Hence, these interventions are as good as those that could be designed by an authority with perfect information.
@@
Furthermore, our results provide tight conditions for robust intervention in the following sense. First, the property of recoverable structure cannot be dispensed with; there exist reasonable demand systems without recoverable structure for which the authority cannot robustly increase total surplus. Second, we show that there are markets with recoverable structure in which any intervention rule that robustly increases total surplus is equivalent, in terms of how it allocates surplus, to the interventions identified by our main result.
@@
We now explain what it means for demand to have a recoverable structure and we discuss how this property is used in the construction of robust interventions.
The demand structure of the oligopoly is reflected in the Slutsky matrix $\bm{D}$. A given cell $D_{ij}$ in this matrix gives us the derivative of product $i$’s demand with respect to the price of product $j$. Thus the matrix specifies the complementarity and substitutability relationships across products. The authority’s signal consists of noisy estimates of the entries of this matrix, with standard assumptions on the errors. The recoverable structure property requires that $\bm{D}$ has some eigenvalues that (in absolute value) grow sufficiently fast as the market gets large, and that the space spanned by their associated eigenvectors correlates with the market demand.
Recoverable structure has both economic and statistical implications that, combined, form the basis of our main result.
The recoverable structure property captures the presence of shared large-scale patterns in the demand of firms. For instance, suppose products can be embedded in a finite-dimensional space, and the location of a product in this space reflects its relation with other products, so that the cell $D_{ij}$ depends on product attributes. If there is a latent set of characteristics relevant for predicting $D_{ij}$ across many products and relevant for predicting market demands (i.e., regressions based on these characteristics have nonvanishing explanatory power) then the matrix $\bm{D}$ would have a recoverable structure.
The statistical aspects of recoverable structure are central to the design of interventions when the Slutsky matrix is observed imperfectly. In particular, the Davis–Kahan Theorem implies that in large markets with a recoverable structure, the noisy observation of $\bm{D}$ can be used to precisely estimate the space of eigenvectors associated with the largest eigenvalues of $\bm{D}$ (Davis and Kahan, 1970).
To use these estimates, we turn to the economic implications of recoverable structure. A key concept in this analysis is pass-through—how changes in firms’ costs affect equilibrium prices, quantities and welfare. We provide a new spectral description of the pass-through of an intervention. That is, we diagonalize the Slutsky matrix to obtain a specific orthonormal basis in which we can express any intervention as a linear combination of orthogonal effects. These effects correspond to the projection of the intervention onto each eigenvector of the Slutsky matrix $\bm{D}$. In large markets with a recoverable structure, an intervention that projects only onto eigenvectors of $\bm{D}$ with large eigenvalues has some very useful properties. First, and most importantly, the orthogonal decomposition means that the welfare effects of such interventions depend only on large eigenvalues of the Slutsky matrix and the subspace they jointly span—the parameters the authority can robustly identify. Second, there are interventions of this form that increase total surplus by more than they cost the authority, providing a net social gain with very low uncertainty. Third, such interventions have minimal effects on prices, which implies that consumer surplus is essentially unchanged, and so the surplus increase achieved by the intervention is incident entirely on producers. Fourth, such interventions actually achieve the maximum possible total surplus per dollar spent subject to the constraint that consumer surplus does not decrease.
@@
Combining the economic and statistical implications of recoverable structure allows us to establish our main result. In markets possessing such structure, the authority can recover precise information about large-eigenvalue components of the spectral decomposition. Using the economic implications of this decomposition that we have just discussed, we then show that interventions based on this information have highly predictable welfare implications, and indeed can be feasibly chosen to achieve the largest possible surplus gain per unit of expenditure subject to not significantly harming consumers.
@@ 1.1. Related literature
Our paper contributes to the literature on the structure and theoretical properties of market power. For an early theoretical paper, see Dixit (1986); more recent studies include, for example, Vives (1999), Azar and Vives (2021), Nocke and Schutz (2018) and Nocke and Whinston (2022). A recent strand of research in macroeconomics and industrial organization uses network models of differentiated oligopoly—similar to the one we consider here—to provide empirical estimates of efficiency losses due to market power (e.g., Pellegrino (2021) and Ederer and Pellegrino (2021)). [^2] See also Elliott and Galeotti (2019) for related arguments about how network methods can be useful for competition authorities in developing antitrust investigations.
@@
Given these estimates of inefficiencies, a natural theoretical question is: What feasible interventions can improve welfare? Our main contribution is to analyze interventions from the perspective of an authority uncertain about the demand structure. Our analysis combines new spectral pass-through formulas with results building on the statistical theory of large matrices, and we identify conditions on the demand structure that ensure the robust achievement of welfare improvements even when many aspects of the demand structure cannot be accurately estimated. [^3] Our focus on pass-through builds on work emphasizing the value of pass-through as a conceptual tool, e.g., Marshall (1890), Pigou (1920), Dixit (1979) and, more recently, Weyl and Fabinger (2013), Miklos-Thal and Shaffer (2021) and Norris (2024). This approach has significant implications for the sorts of empirical models that are needed to design interventions in large markets with many goods. We elaborate on these issues in Section 7.2.
@@
Methods in high-dimensional statistics are currently attracting considerable interest in econometric settings where covariates are high-dimensional (see, e.g., Athey, Bayati, Doudchenko, Imbens, and Khosravi (2021) and Chernozhukov, Hansen, Liao, and Zhu (2023)), and there is work applying related statistical models to informational or behavioral spillovers in social networks (Golub and Jackson, 2012; Dasaratha, 2020; Cai, 2022; Parise and Ozdaglar, 2023; Chandrasekhar, Goldsmith-Pinkham, McCormick, Thau, and Wei, 2024). However, we know little about the question of when noisy data can be effectively used in order to implement desirable interventions in the presence of strategic spillovers, particularly when it comes to market settings. We show that, in a large oligopoly market, the concepts developed in the literature on large network recovery can be useful for the design of socially desirable interventions.
@@
Finally, our paper contributes to the theory of network interventions. Early contributions in this field include Borgatti (2006), Ballester, Calvó-Armengol, and Zenou (2006), and Goyal (1996). [^4] The literature on this subject is very large. Intervention design has been studied in models of information diffusion, advertising, finance, security, and pricing, among other topics—see e.g., Banerjee, Chandrasekhar, Duflo, and Jackson (2013), Bloch and Querou (2013), Candogan, Bimpikis, and Ozdaglar (2012), Elliott et al. (2014), Belhaj and Deroian (2017), Demange (2017), Dziubinski and Goyal (2017), Galeotti and Goyal (2009), Galeotti and Rogers (2013), and Leduc, Jackson, and Johari (2017). In recent years, spectral methods have been applied to study optimal intervention problems under the assumption that the authority has perfect information about the network of spillovers (Galeotti, Golub, and Goyal, 2020; Gaitonde, Kleinberg, and Tardos, 2021; Liu and Tsyvinski, 2024). [^5] Some recent work uses spectral analysis to derive conditions for core-selecting re-allocative auctions (Rostek and Yoder, 2023), and robust implementation (Ollár and Penta, 2023). By contrast, in the present paper, we study a setting in which the authority observes the structure of strategic spillovers with significant noise. [^6] We share with several prior papers the idea that decision-makers act under partial information about the network. For instance, Galeotti, Goyal, Jackson, Vega-Redondo, and Yariv (2010) study large network games where players have incomplete information about the network structure, described by a random graph; Akbarpour, Malladi, and Saberi (2020) considers seeding in a large random graph; the diffusion process there lacks any form of complementarity. Our questions and methods of analysis are very different from these papers. The methods we develop for robust interventions can be applied to other network games more generally and we briefly discuss this in Section 7.3.
2. Model
In this section we describe the differentiated oligopoly game, the interventions available to the authority, and our notion of robust interventions.
@@ 2.1.1. Demand side
There is a set $\{1,...,n\}$ of distinct products. The demand for these products arises from the consumption choices of many optimizing households. Each household $h\in\{1,\dots,H\}$ takes prices as given and has a choice utility that is quasilinear in a numeraire $m$,
@@
where $V^{h}$ is a twice-differentiable and strictly concave function of the consumption profile $\tilde{\bm{q}}^{h}\in\mathbb{R}^{n}$ and $m$ is a numeraire (“money”), in which all prices are denominated. Given a price profile $\tilde{\bm{p}}$, the household’s problem is to choose a bundle $\tilde{\bm{q}}^{h}$ to maximize ${U}^{h}(\tilde{\bm{q}}^{h},m)-\tilde{\bm{p}}\cdot\tilde{\bm{q}}^{h}$. Letting $\bm{q}^{h}(\tilde{\bm{p}})$ be the (unique, by concavity of $V^{h}$) solution to household $h$’s problem, total market demand is777 Because the households’ utilities are quasilinear in money, one can derive the same aggregate demand from a representative consumer, and this description is sufficient for studying producer and aggregate consumer surplus. However, some of our results will provide more refined results about the effects of consumers, showing that no household is significantly hurt.

### v2

@@
When can interventions in markets be designed to increase surplus robustly—i.e., with high probability—accounting for uncertainty due to imprecise information about economic primitives? In a setting with many strategic firms, each possessing some market power, we present conditions for such interventions to exist. The key condition, recoverable structure, requires large-scale complementarities among families of products. The analysis works by decomposing the incidence of interventions in terms of principal components of a Slutsky matrix. Under recoverable structure, a noisy signal of this matrix reveals enough about these principal components to design robust interventions. Our results demonstrate the usefulness of spectral methods for analyzing imperfectly observed strategic interactions with many agents.
Robust Market Interventions
Andrea Galeotti and Benjamin Golub and Sanjeev Goyal and
Eduard Talamàs and Omer Tamuz
(Date Printed. December 31, 2024)
Galeotti is at the London Business School, Golub is at Northwestern University, Goyal is at University of Cambridge and NYUAD, Talamàs is at IESE Business School, and Tamuz is at Caltech.
       We thank Yann Calvó López and Yu-Chi Hsieh for excellent research assistance.
      This work was supported by the ERC grant 724356 (Galeotti), the Pershing Square Fund for Research on the Foundations of Human Behavior (Golub), the National Science Foundation (SES-1658940, SES-1629446, Golub & DMS-1944153, Tamuz), the Sloan Foundation (Tamuz) and the US-Israel Binational Science Foundation (2018397, Tamuz).
       We thank Bruno Pelligrino for valuable exchanges and kindly sharing data with us, and Volker Nocke for a helpful discussion on a related project at the 2023 AEA meetings. We thank Pol Antràs, Itai Ashlagi, Modibo Camara, Yeon-Koo Che, Kevin Chen, Rebecca Diamond, Jeff Ely, Joey Feffer, Alex Frankel, Matthew Gentzkow, Tim Gowers, Matthew O. Jackson, Zi Yang Kang, Ernest Liu, Margaret Meyer, Jose Luis Moraga, Stephen Morris, Martin Pietz, Ilya Segal, Andy Skrzypacz, Ran Spiegler, Ludwig Straub, Xavier Vives, Kuang Xu, and Frank Yang for helpful comments.
       The paper subsumes the working paper Galeotti et al. (2022).
@@ 1. Introduction
Market power has recently attracted renewed attention and is thought to have significant and growing welfare implications (see, e.g., Syverson, 2019). While many applied analyses of competition confine the analysis of market power to tightly defined product markets, it is becoming clear that important welfare-relevant spillovers operate across such markets (Baqaee and Farhi, 2020; Azar and Vives, 2021; Pellegrino, 2021; Ederer and Pellegrino, 2021). Our theoretical understanding of such spillovers in environments with many firms interacting via general demand systems remains limited. This paper is about the welfare theory of such interactions.
Consider many profit-maximizing, single-product firms simultaneously setting prices, with arbitrary complementarities and substitutabilities across products. For instance, one firm’s product—e.g., a Samsung smartphone—may be a substitute to some products—e.g., Apple smartphones—and a complement to others—compatible accessories such as earbuds, watches, and smart home appliances, which may in turn be substitutes or complements to one another.
We are interested in the nature of inefficiencies in such an environment and policies to respond to them. We consider these issues from the perspective of an authority that recognizes the possibility of inefficiency due to market power and can intervene through taxes and subsidies on firms’ sales. For concreteness, we will think of this authority as the operator of a large marketplace, such as Amazon, that mediates retail sales and aims to increase the equilibrium economic surplus generated by the marketplace. [^1] The aim of this might be to deliver more value to consumers and producers to keep the platform competitive, or to reclaim the surplus via lump-sum fixed fees. What principles should guide the design of such interventions? When can such policies be implemented under realistic uncertainty about market parameters?
What makes the problem challenging is that, once we broaden our perspective beyond one traditionally defined market (e.g., smartphones) and consider spillovers to a variety of other complements and substitutes, there is a kind of curse of dimensionality. In a marketplace with numerous and changing goods, the demand system is high-dimensional and a priori unstructured. As any firm’s cost changes, the number of potential effects to consider is equal to the number of products; thus, the number of interactions quadratically in this number. Realistic signals will leave substantial uncertainty about many aspects of the structure of the game among the firms (see Section 7.1 for a detailed discussion). In particular, the authority will have nothing close to a precise estimate of the entire demand system. This raises the question is whether there are policies that robustly improve surplus despite this uncertainty.
@@
Our main result is that if demand satisfies a property that we call recoverable structure, then there are feasible intervention rules that robustly (i.e., with high probability) increase equilibrium total surplus despite large errors in observing every detail of the system. Moreover, within a natural class of interventions—those that do not reduce consumer surplus [^2] This constraint is natural for a platform that could lose customers to other marketplaces or a government that can lose political support.—our feasible interventions achieve the largest gain in surplus that is possible for a given level of subsidy expenditure. Hence, within this class, these interventions are as good as those that could be designed by an authority with perfect information.
@@
Furthermore, our results provide tight conditions for robust intervention in the following sense. First, the property of recoverable structure cannot be dispensed with; there are reasonable demand systems without recoverable structure for which the authority cannot robustly increase total surplus. Second, we show that there are settings with recoverable structure in which any intervention rule that robustly increases total surplus is equivalent, in terms of how it allocates surplus, to the interventions identified by our main result. These tightness results show that robust interventions must, in general, be tailored to the marketplace—there is no simple rule of thumb that always works. On the other hand, under conditions that we identify, it is possible to tailor rules well, despite large uncertainty in many aspects of demand.
@@
The key condition in the paper is recoverable structure. We now explain what it means for demand to have recoverable structure and how this property is used in the construction of robust interventions.
The demand structure is encoded in a matrix $\bm{D}$ of demand derivatives, which in our setting is equal to the Slutsky matrix. A given cell $D_{ij}$ in this matrix is the derivative of product $i$’s demand with respect to product $j$’s price. Thus, the matrix specifies the complementarity and substitutability relationships across products. Mathematically, the recoverable structure property requires that $\bm{D}$ can be written as a rank-one matrix with large norm plus a matrix orthogonal to this. This rank-one piece can be thought of as a large principal component: a part of the demand system described by a single vector that accounts for a large amount of demand behavior. In terms of the economic intuition, we will show that recoverable structure entails substantial large-scale complementarities. This manifests as the ability of small subsidies to have large spillover effects that raise the consumption of many goods by significant amounts.
The key point here is that recoverable structure is large-scale structure. When large-scale complementarities are present, they correspond to marketplace-wide double marginalization problems, where many goods all exert externalities on one another. Such broad externalities create the potential for interventions that substantially improve welfare.
The statistical implications of recoverable structure are then central to actually taking advantage of this potential when the Slutsky matrix is observed imperfectly. The authority’s signal consists of noisy estimates of the entries of this matrix, with noise magnitudes in each entry comparable to the entries themselves. This noise creates large uncertainty in the operation of a given intervention. We show that, nevertheless, in large markets with recoverable structure, such noisy observation of $\bm{D}$ can be used to precisely predict the effects of some well-chosen interventions—specifically, those operating in the space of eigenvectors associated with the largest eigenvalues of $\bm{D}$. The key tool for this is the Davis–Kahan theorem (Davis and Kahan, 1970).
@@
Combining the economic and statistical implications of recoverable structure allows us to establish our main result. In markets possessing such structure, the authority can recover precise information about large-eigenvalue components of the spectral decomposition, and interventions based on this information have highly predictable surplus implications.
At a technical level, to perform this analysis we develop a new spectral description of the pass-through of an intervention. That is, we diagonalize the Slutsky matrix to obtain a specific orthonormal basis in which we can express the implications of any intervention as a linear combination of orthogonal effects. These effects correspond to the projection of the intervention onto each eigenvector of the Slutsky matrix $\bm{D}$. By characterizing the pass-throughs of subsidies to prices, quantities, and welfare separately across these principal components, we are able to prove that targeting the high-eigenvalue principal components yields precisely predictable results achieving our claimed welfare properties. The spectral decomposition may be of independent interest, yielding a useful basis in which price and welfare pass-throughs of cost shocks behave intuitively despite the complexity of a system with arbitrary spillovers.
@@ 1.1. Related literature
Our paper contributes to the literature on the structure and theoretical properties of market power. For an early theoretical paper, see Dixit (1986); more recent studies include, for example, Vives (1999), Azar and Vives (2021), Nocke and Schutz (2018), and Nocke and Whinston (2022). A recent strand of research in macroeconomics and industrial organization uses differentiated oligopoly network models—similar to the one we consider here—to provide empirical estimates of efficiency losses due to market power (e.g., Pellegrino (2021) and Ederer and Pellegrino (2021)). [^3] See also Elliott and Galeotti (2019) for related arguments about how network methods can be useful for competition authorities in developing antitrust investigations.
@@
Given these estimates of inefficiencies, a natural theoretical question is: What feasible interventions can improve welfare? Our main contribution is to analyze interventions from the perspective of an authority uncertain about the demand structure. Our analysis combines new spectral pass-through formulas with results building on the statistical theory of large matrices, and we identify conditions on the demand structure that ensure the robust achievement of welfare improvements even when many aspects of the demand structure cannot be accurately estimated. [^4] Our focus on pass-through builds on work emphasizing the value of pass-through as a conceptual tool, e.g., Marshall (1890), Pigou (1920), Dixit (1979) and, more recently, Weyl and Fabinger (2013), Miklos-Thal and Shaffer (2021) and Norris (2024). This approach has significant implications for understanding which kinds of empirical models are needed to design interventions in large markets with many goods. We elaborate on these issues in Section 7.2.
@@
Methods in high-dimensional statistics are currently attracting considerable interest in econometric settings with high-dimensional covariates (see, e.g., Athey, Bayati, Doudchenko, Imbens, and Khosravi (2021) and Chernozhukov, Hansen, Liao, and Zhu (2023)), and there is work applying related statistical models to informational or behavioral spillovers in social networks (Golub and Jackson, 2012; Dasaratha, 2020; Cai, 2022; Parise and Ozdaglar, 2023; Chandrasekhar, Goldsmith-Pinkham, McCormick, Thau, and Wei, 2024). However, we know little about when noisy data can be effectively used in order to implement desirable interventions in the presence of strategic spillovers, particularly in market settings. We show that, in a large oligopoly market, the concepts developed in the literature on large network recovery can be useful for designing socially desirable interventions.
@@
Our paper contributes to the theory of network interventions. Early contributions include Borgatti (2006), Ballester, Calvó-Armengol, and Zenou (2006), and Goyal (1996). [^5] The literature on this subject is very large. Intervention design has been studied in models of information diffusion, advertising, finance, security, and pricing, among other topics—see e.g., Banerjee, Chandrasekhar, Duflo, and Jackson (2013), Bloch and Querou (2013), Candogan, Bimpikis, and Ozdaglar (2012), Belhaj and Deroian (2017), Demange (2017), Dziubinski and Goyal (2017), Galeotti and Goyal (2009), and Leduc, Jackson, and Johari (2017). Spectral methods have recently been applied to optimal intervention problems when spillovers are known (Galeotti, Golub, and Goyal, 2020; Gaitonde, Kleinberg, and Tardos, 2021; Liu and Tsyvinski, 2024). [^6] Some recent work uses spectral analysis to derive conditions for core-selecting re-allocative auctions (Rostek and Yoder, 2023), and robust implementation (Ollár and Penta, 2023). See also Aguiar and Serrano (2017) on spectral methods to study Slutsky matrices in a consumer theory setting. By contrast, in the present paper, the authority observes strategic spillovers with significant noise. [^7] We share with several prior papers the idea that decision-makers act under partial information about the network. For instance, Galeotti, Goyal, Jackson, Vega-Redondo, and Yariv (2010) study large network games where players have incomplete information about the network structure, described by a random graph; Akbarpour, Malladi, and Saberi (2020) considers seeding in a large random graph; the diffusion process there lacks any form of complementarity. Our questions and methods of analysis are very different from these papers. The methods we develop for robust interventions can be applied to other network games more generally and we briefly discuss this in Section 7.3. Our analysis of perturbations of taxes and subsidies is related to the classic “tax reform approach” in public finance (Feldstein, 1976; Tirole and Guesnerie, 1981); the study of uncertain spillovers distinguishes our work.
Our approach to robustness is conceptually related to, but methodologically distinct from, an extensive literature in economic theory. That literature focuses on understanding the design of mechanisms and contracts that achieve desired outcomes even when assumptions about the environment (e.g., agents’ preferences, beliefs, and rationality) are relaxed; see Carroll (2019) for a survey. Our definition of robustness aligns with the spirit of this literature. However, in our context, the motivation for analyzing robust interventions arises from the high-dimensional nature of the market state, and use methods that align with statistical work in this type of setting.
2. Framework
In this section, we present the framework for our study. The foundation is a simple differentiated oligopoly game. Within this game, we introduce a class of interventions available to the authority and calculate the surplus outcomes associated with these interventions. Finally, we introduce the statistical framework describing the signals available to the authority and a notion of rules that use these signals to achieve good outcomes robustly.
@@ 2.1.1. Demand side
There is a set $\{1,\ldots,n\}$ of distinct products. The demand for these products arises from the consumption choices of a fixed, finite number of optimizing households. Each household $h\in\{1,\dots,H\}$ takes prices as given and has a choice utility that is quasilinear in a numeraire $m$,
@@
where $V^{h}$ is a twice-differentiable and strictly concave function of the consumption profile $\tilde{\bm{q}}^{h}\in\mathbb{R}^{n}$ and $m$ is a numeraire (“money”), in which all prices are denominated. Given a price profile $\tilde{\bm{p}}$, the household’s problem is to choose a bundle $\tilde{\bm{q}}^{h}$ to maximize ${U}^{h}(\tilde{\bm{q}}^{h},m)-\tilde{\bm{p}}\cdot\tilde{\bm{q}}^{h}$. Letting $\bm{q}^{h}(\tilde{\bm{p}})$ be the solution to household $h$’s problem (unique by concavity of $V^{h}$), total market demand is888 Because the households’ utilities are quasilinear in money, one can derive the same aggregate demand from a representative consumer, and this description is sufficient for studying producer and aggregate consumer surplus. However, some of our results will provide more refined results about the effects on individual consumers, showing that no household is significantly hurt.